# Is LLM-as-a-Judge Robust? Investigating Universal Adversarial Attacks on Zero-shot LLM Assessment

### v1

@@ Abstract
Large Language Models (LLMs) are powerful zero-shot assessors and are increasingly used in real-world situations such as for written exams or benchmarking systems. Despite this, no existing work has analyzed the vulnerability of judge-LLMs against adversaries attempting to manipulate outputs. This work presents the first study on the adversarial robustness of assessment LLMs, where we search for short universal phrases that when appended to texts can deceive LLMs to provide high assessment scores. Experiments on SummEval and TopicalChat demonstrate that both LLM-scoring and pairwise LLM-comparative assessment are vulnerable to simple concatenation attacks, where in particular LLM-scoring is very susceptible and can yield maximum assessment scores irrespective of the input text quality. Interestingly, such attacks are transferable and phrases learned on smaller open-source LLMs can be applied to larger closed-source models, such as GPT3.5. This highlights the pervasive nature of the adversarial vulnerabilities across different judge-LLM sizes, families and methods. Our findings raise significant concerns on the reliability of LLMs-as-a-judge methods, and underscore the importance of addressing vulnerabilities in LLM assessment methods before deployment in high-stakes real-world scenarios.  [^1] Code: https://github.com/rainavyas/attack-comparative-assessment
@@
Vyas ${}^{\ast}$ University of Cambridge vr313@cam.ac.uk                        Adian ${}^{\ast}$ University of Cambridge al826@cam.ac.uk                        Mark Gales University of Cambridge mjfg@cam.ac.uk
${}^{\ast}$ Equal Contribution.
@@ 1 Introduction
This work is the first to apply adversarial attacks (Szegedy et al., 2014) to zero-shot LLM assessment, and demonstrates that LLM-as-a-judge methods are susceptible to simple concatenative adversarial attacks. Both LLM-scoring and pairwise LLM-comparative assessment are vulnerable to such attacks, and concatenating a universal phrase of under 5 tokens can trick systems into providing significantly higher assessment scores. We find that comparative assessment has mild robustness, but that LLM-scoring is very vulnerable to such attacks and universal attack phrases can cause systems to return a maximum score, irrespective of the input text. We further demonstrate that for LLM-scoring, phrases can be transferable across different model sizes and families, where attack phrases learned on FlanT5-3B can deceive much more powerful systems such as ChatGPT. Finally, as an initial step towards defending against such attacks, we use the perplexity score (Jain et al., 2023) as a simple detection approach, which demonstrates some success. As a whole, our work raises awareness of the vulnerabilities of zero-shot LLM assessment, and highlights that if such systems are to be deployed in critical real-world scenarios, such vulnerabilities should be considered and addressed.
@@ 3 Zero-shot Assessment with LLMs
  • LLM Absolute Scoring where an LLM is asked to assign an absolute score to each considered text.
@@ 3.1 Comparative Assessment
3.2 Absolute Assessment
In LLM absolute scoring, the LLM, $\mathcal{F}$, is prompted to directly predict an absolute score. The prompt is designed to request the LLM to assess the quality of a text with a score (e.g. between 1-5). Two variants of absolute assessment can be applied; first where the score is directly predicted by the LLM,
@@ 4 Adversarial Assessment Attacks
4.1 Attack Objective
@@ 4.2 Practical Attack Approach
In this work, we use a simple greedy search to learn the universal attack phrase  [^2] We also carried out experiments using the Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023) to learn the universal attack phrase, but this approach was found to be not as effective as the greedy search process. Results for GCG experiments are provided in Appendix C.. For a vocabulary, $\mathcal{V}$ the greedy search finds the most effective adversarial word to append iteratively,
@@
Algorithm 1 Greedy Search Universal Attack for Comparative Assessment
@@
              $s_{a}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a}+\bm{\delta},\mathbf{x}^{(m)}_{b},\mathbf{c}^{(m)})$
              $s_{b}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a},\mathbf{x}^{(m)}_{b}+\bm{\delta},\mathbf{c}^{(m)})$
              $q\leftarrow q+s_{a}+s_{b}$
         end for
         if $q>q^{*}$ then
              $q^{*}\leftarrow q$
              $\delta^{*}_{l}\leftarrow\delta$triangleright Update best attack word
         end if
     end for
end for
Algorithm 2 Greedy Search Universal Attack for Absolute Assessment
$\mathcal{F}()$triangleright Target Model
$\bm{\delta}^{*}\leftarrow\text{ empty string}$
for $l=1:L$ do
     $a\sim\{1,...,N\}$triangleright Select candidate index
     $\delta^{*}_{l}\leftarrow\text{ none}$
     $q^{*}\leftarrow 0$triangleright Initialize best score
     for $\delta\in\mathcal{V}$ do
         $q\leftarrow 0$
         for $m=1:M$ do
@@ 5.2 LLM Assessment Systems
We investigate a range of standard instruction-tuned generative language models in our experiments; FlanT5-xl (3B parameters) Chung et al. (2022); Llama2-7B-chat Touvron et al. (2023); Mistral-7B-chat Jiang et al. (2023) and GPT3.5 (175B parameters). For both comparative and absolute assessment, we perform all attacks on FlanT5-xl. Our prompts for comparative assessment follow the prompts used in Liusie et al. (2023), where different attributes use different adjectives in the prompt. For absolute assessment we follow the prompts of G-Eval Liu et al. (2023b) and use continuous scores (Equation 4) by calculating the expected score over a score range (e.g. 1-5 normalized by their probabilities). Note that the GPT3.5 API does not give token probabilities, and so for GPT3.5 we use standard prompt scoring (Equation 3).
@@ 5.3 Methodology
Each dataset is split into a development set and a test set following a 20:80 ratio. We use the development set (20% of the passages) to learn the attack phrase using the simple greedy search to maximise the expected score of the attacked sample (§4.2) and evaluate using the test set (80% of the passages). Furthermore, we only use two of the candidate texts to learn the attacks (i.e. 2 of 16 for SummEval and 2 of 6 for topical chat), and therefore perform the search over a modest 40 summaries for SummEval and 24 responses for TopicalChat.
We learn a single universal attack phrase on the target model, which we keep as FlanT5-xl for all learned phrases. We perform a separate universal concatenation attack for each dataset and attribute, and use the notation ({TASK} {ASSESSMENT} {ATTRIBUTE}) to indicate the task, the assessment type and the evaluation attribute for each learnt universal attack phrase. We specifically consider the following configurations: (SUMM COMP OVE); (TOPIC COMP OVE); (SUMM ABS OVE); (TOPIC ABS OVE); (SUMM COMP CON); (TOPIC COMP CNT); (SUMM ABS CON); (TOPIC ABS CNT)  [^3] The learnt universal attack phrases for each configuration are given in Appendix A.. The specific attributes CON and CNT are selected due to the smallest performance difference in comparative and absolute assessment (Tables 1 and 2). The attack phrases are then assessed on transferability to the other target models; Mistral-7B, Lllama2-7B, GPT3.5. The vocabulary for the greedy attack is sourced from the NLTK python package  [^4] English words corpus is sourced from: nltk.corpus.
@@ 6 Results
Figure 2: Universal Attack Evaluation (average rank of attacked summary/response) for FlanT5-xl.
@@ 6.1 Assessment Performance
Tables 1 and 2 present the assessment ability of each LLM when applied to comparative and absolute assessment for SummEval and TopicalChat. Consistent with literature, comparative assessment performs better than absolute assessment systems for most systems and attributes. However, this comes with greater computational complexity, as $N\!\cdot\!(N-1)$ comparisons are required to compare all pairs of responses (Equation 2), whilst only $N$ inferences are required for absolute assessment. Smaller LLMs (FlanT5-xl, Llama2-7b and Mistral-7b) demonstrate reasonable performance on SummEval and TopicalChat, but larger models (GPT3.5) perform much better, and even when using simple absolute assessment, can outperform smaller models applying comparative assessment.
6.2 Universal Adversarial Attacks
Section 5.3 details the attack configuration and approach used to learn the universal attack phrases. Figure 2 illustrates the impact the universal adversarial attack phrase can have for comparative and absolute assessments on SummEval and TopicalChat evaluation while using FlanT5-xl as the LLM assessment system. For Summeval we attack the overall score (OVE) and the consistency (CON) while for Topical-Chat we attack the overall score (OVE) and the continuity (CNT). The attributes CON and CNT were selected due to the similar observed performance of the absolute and comparative methods on these attributes.
The success of the adversarial attacks is measured by the average ranks of the text after an attack. Figure 2 demonstrates that both comparative assessment and absolute assessment systems have some vulnerability to adversarial attacks, as the average rank decreases, and continues to decrease as more words are added to the attack phrase. However, absolute assessment systems are significantly more susceptible to universal adversarial attacks, and with just four universal attack words, the absolute assessment system will consistently provide a rank of 1 to nearly all input texts. Table 3 provides the raw scores for comparative and absolute assessment, where we see that for absolute assessment, a universal attack phrase of 4 words will yield assessment scores on average near the maximum score of 5. The specific universal attack phrases learnt for each task are given in Appendix A.
@@ 6.2 Universal Adversarial Attacks
Figure 3: Transferability of universal attack phrases from FlanT5-xl to other models.
The relative robustness of comparative assessment systems over absolute assessment systems can perhaps be explained intuitively. In an absolute assessment setting, an adversary exploits an input space which is not well understood by the model and identifies a region that spuriously encourages the model to predict a high score. However, in comparative assessment, the model is forced to compare the quality of the attacked text to another (unattacked) text, meaning the attack phrase learnt has to be invariant to the text used for comparison. This makes it more challenging to find an effective adversarial attack phrase. Further explanations for the relative robustness of comparative assessment systems are explored in Appendix B.
6.3 Transferability
Figure 2 demonstrated that absolute assessment systems are highly vulnerable to a simple universal attack phrase concatenated to an input text. However, running an adversarial attack on black-box systems may not be feasible due to the large number of calls required for a greedy search, assuming a reasonable vocabulary of $|\mathcal{V}|\approx 50,000$. To circumvent this issue, an adversary can learn the universal attack-phrase on a smaller open-source model (with no detectable limit on the number of model queries) and aim to transfer this attack phrase to other models (e.g., in Demontis et al. (2018)).
The previous universal attack phrases of Figure 2 was learned specifically for FlanT5-xl, which is a relatively poor-performing model. However, Flan-T5 is quite small (3B parameters); one may therefore consider whether attack phrases learned using this smaller model can generalize and deceive larger models. Figure 3 shows that 1) there can be a high level of attack transferability for absolute scoring. For TopicalChat, it appears that the attacks generalize very well to nearly all systems, with all systems being very susceptible to attacks when assessing continuity. 2) It appears that when more powerful models assess the overall (OVE) quality, there can be less effective transferability, suggesting that assessing more general, abstract qualities can be more robust. Further, interestingly it appears that powerful large models (GPT3.5) are more susceptible when attacked by shorter phrases, possibly as the longer phrase may begin to overfit to properties of the attacked model. 3) The attack transfers with mixed success for Summeval, which may highlight that the complexity of dataset can influence attack transferability.
@@ 6.4 Attack Detection
In this section, we perform an initial investigation into possible defences that could be applied to detect if an adversary is exploiting a system. Defences can take two forms: adversarial training (Goodfellow et al., 2015) where the LLM is re-trained with adversarial examples, or adversarial attack detection where a separate module is designed to identify adversarial inputs. Although recent LLM adversarial training approaches have been proposed (Zhou et al., 2024; Zhang et al., 2023b), re-training is computationally expensive and can harm model performance, hence detection is preferred. Recent detection approaches for NLG adversarial attacks tend to focus on attacks that circumvent LLM safety filters, e.g., generateing malicious content by jailbreaking (Liu et al., 2023c; Zou et al., 2023; Jin et al., 2024). Robey et al. (2023) propose SmoothLLM, where multiple versions of the perturbed input are passed to an LLM and the outputs aggregated. Such defences are inappropriate for LLM-as-a-judge setups, as though the perturbations are designed to cause no semantic change, they can result in changes in other attributes, such as fluency and style, which will impact the LLM assessment. Similarly, Jain et al. (2023); Kumar et al. (2024) propose defence approaches that involve some form of paraphrasing or filtering of the input sequence, which again interferes with the LLM-as-a-judge scores.
@@
$\texttt{P}=\frac{\text{TP}}{\text{TP+FP}}\quad\texttt{R}=\frac{\text{TP}}{\text{TP+FN}}\quad F1=2\cdot\frac{\texttt{P}\cdot\texttt{R}}{\texttt{P}+\texttt{R}},$
@@ 7 Conclusions
This is the first work to examine the adversarial robustness of zero-shot LLM assessment methods against universal concatenation adversarial attacks. Our findings reveal vulnerabilities in both LLM comparative assessment and LLM scoring, with absolute assessment exhibiting particularly high susceptibility. Notably, we demonstrate that attacks devised on smaller systems can effectively transfer to larger ones, underscoring the need for defence strategies. Encouragingly, an initial investigation into detection defences, shows that perplexity can be a promising tool for identifying adversarially manipulated inputs. On the whole, this paper raises awareness around the susceptibility of LLM-as-a-judge NLG assessment systems to universal and transferable adversarial attacks. Further work can explore adaptive attacks and more sophisticated defence approaches to minimize the risk of misuse.

### v2

@@ Abstract
Large Language Models (LLMs) are powerful zero-shot assessors used in real-world situations such as assessing written exams and benchmarking systems. Despite these critical applications, no existing work has analyzed the vulnerability of judge-LLMs to adversarial manipulation. This work presents the first study on the adversarial robustness of assessment LLMs, where we demonstrate that short universal adversarial phrases can be concatenated to deceive judge LLMs to predict inflated scores. Since adversaries may not know or have access to the judge-LLMs, we propose a simple surrogate attack where a surrogate model is first attacked, and the learned attack phrase then transferred to unknown judge-LLMs. We propose a practical algorithm to determine the short universal attack phrases and demonstrate that when transferred to unseen models, scores can be drastically inflated such that irrespective of the assessed text, maximum scores are predicted. It is found that judge-LLMs are significantly more susceptible to these adversarial attacks when used for absolute scoring, as opposed to comparative assessment. Our findings raise concerns on the reliability of LLM-as-a-judge methods, and emphasize the importance of addressing vulnerabilities in LLM assessment methods before deployment in high-stakes real-world scenarios.  [^1] Code: https://github.com/rainavyas/attack-comparative-assessment
@@
Vyas Raina University of Cambridge vr313@cam.ac.uk                        Adian Liusie University of Cambridge al826@cam.ac.uk                        Mark Gales University of Cambridge mjfg@cam.ac.uk
Equal Contribution.
@@ 1 Introduction
This work is the first to propose adversarial attacks Szegedy et al. (2014) targeting zero-shot LLM assessment. In practical settings, the adversary may either not have any knowledge of the judge-LLMs, access to the model weights, or be limited in the number of queries that can be made to the model (due to costs or suspicion from excessive querying). Therefore, we learn the attack phrase while using a surrogate model  (Papernot et al., 2016) and transfer the universal attack phrase to other judge-LLMs. We demonstrate that universal attack phrases learned with access only to FlanT5-3B model, a small encoder-decoder transformer, can transfer to larger decoder-only models and cause Llama2-7B, Mistral-7B and ChatGPT to return the maximum score, irrespective of the input text. We find that LLM-scoring (as opposed to pairwise LLM-comparative assessment) can be particularly vulnerable to such attacks, and concatenating a universal phrase of just 5 tokens can trick these systems into providing highly increased assessment scores. Additionally, we find that comparative assessment is more robust than LLM-scoring to such adversarial attacks, although the direct attacks on the surrogate model can yield marginally inflated scores. Finally, as an initial step towards defending against such attacks, we use the perplexity score (Jain et al., 2023) as a simple detection approach, which demonstrates some success. As a whole, our work raises awareness of the vulnerabilities of zero-shot LLM assessment, and highlights that if such systems are to be deployed in critical real-world scenarios, adversarial vulnerabilities should be considered and addressed.
@@ 3 Zero-shot Assessment with LLMs
  • LLM Scoring where an LLM is asked to assign an absolute score to each considered text.
@@ 3.1 Comparative Assessment
3.2 Absolute Scoring Assessment
In LLM absolute scoring, the LLM, $\mathcal{F}$, is prompted to directly predict the assessment score. The prompt is designed to request the LLM to assess the quality of a text with a score (e.g. between 1-5). Two variants of scoring can be applied; first where the score is directly predicted by the LLM,
@@ 4 Adversarial Assessment Attacks
4.1 Attack Threat Model
Objective.
@@ 4.1 Attack Threat Model
Surrogate Model Transfer Attack.
Traditional adversarial attack methods often assume full access to the target model, but this setting might be unrealistic when attacking assessment systems. Hence, we consider the more practical scenario where the adversary only has full access to a surrogate model that differs from the actual judge-LLM used by the assessment system. The attack can be learned on the surrogate model and then transferred to the target model as initially proposed by Liu et al. (2016); Papernot et al. (2016). The assumption is that due to possible similarities in training data, training recipes and model architectures, the attacks may transfer reasonably to the target model.
@@ 4.2 Practical Attack Approach
In this work, we use a simple greedy search to learn the universal attack phrase  [^2] We also carried out experiments using the Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023) to learn the universal attack phrase, but this approach was found to be not as effective as the greedy search process. Results for GCG experiments are provided in Appendix E.. For a vocabulary, $\mathcal{V}$ the greedy search finds the most effective adversarial word to append iteratively,
@@
The algorithm for the practical greedy search attack on comparative assessment and absolute assessment systems is given in Algorithm 1.
Algorithm 1 Greedy Search Universal Attack for LLM Comparative Assessment LLM and Scoring
@@
              if comparative then
                  $p_{1}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a}+\bm{\delta},\mathbf{x}^{(m)}_{b},\mathbf{c}^{(m)})$
                  $p_{2}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a},\mathbf{x}^{(m)}_{b}+\bm{\delta},\mathbf{c}^{(m)})$
                  $q\leftarrow q+p_{1}+(1\!-\!p_{2})$
              else if scoring then
@@ 5.2 LLM Assessment Systems
We consider a range of standard instruction-tuned generative language models that can be used as judge-LLMs: FlanT5-xl (3B parameters) Chung et al. (2022), Llama2-7B-chat Touvron et al. (2023), Mistral-7B-chat Jiang et al. (2023), and GPT3.5 (175B parameters). FlanT5-xl, the smallest and the only encoder-decoder system, is used as the surrogate model for learning the universal adversarial attack phrases for both comparative and absolute assessment. Once the attack phrases are learned on FlanT5-xl, they are transferred to the other target LLMs to evaluate their effectiveness. Our prompts for comparative assessment follow the prompts used in Liusie et al. (2023), where different attributes use different adjectives in the prompt. For absolute assessment, we follow the prompts of G-Eval Liu et al. (2023b) and use continuous scores (Equation 4) by calculating the expected score over a score range (e.g., 1-5 normalized by their probabilities). Note that the GPT3.5 API does not provide token probabilities, so for GPT3.5, we use standard prompts without token probability normalization.
@@ 5.3 Methodology
Each dataset is split into a development set and a test set following a 20:80 ratio. We use the development set (20% of the passages) to learn the attack phrase using a simple greedy search to maximize the expected score of the attacked samples and evaluate using the test set (80% of the passages). Furthermore, we only use two of the candidate texts to learn the attacks (i.e., 2 of 16 for SummEval and 2 of 6 for TopicalChat), and therefore perform the search over a modest total of 40 summaries for SummEval and 24 responses for TopicalChat.
For each dataset and attribute, we perform a separate universal concatenation attack using the notation (TASK ASSESSMENT ATTRIBUTE) to indicate the task (SummEval, TopicalChat), the assessment method (comparative, scoring), and the evaluation attribute (overall, consistency, continuity) for each learned universal attack phrase  [^3] The learned universal attack phrases for each configuration are given in Appendix A.. E.g., SUMM-COMP-OVE denotes the phrase learned for comparative assessment when attacking the SummEval overall score.
We learn a single universal attack phrase on the surrogate model, FlanT5-xl, for all experiments in the main paper. Once the universal attack phrases are learned on the surrogate model, the attack is further assessed when transferred to the other target models: Mistral-7B, Llama2-7B, and GPT3.5. The vocabulary for the greedy attack is sourced from the NLTK python package  [^4] English words corpus is sourced from: nltk.corpus.
@@ 6 Results
Figure 2: Universal attack evaluation (average rank of attacked summary/response) for surrogate FlanT5-xl.
@@ 6.1 Assessment Performance
Tables 1 and 2 present the assessment ability of each LLM when applied to comparative and absolute assessment for SummEval and TopicalChat. Consistent with literature, comparative assessment performs better than absolute assessment systems for most systems and attributes. However, comparative assessment uses $N\!\cdot\!(N\!-\!1)$ to compare all pairs of responses (Equation 2), whilst only $N$ inferences are required for absolute assessment. Smaller LLMs (FlanT5-xl, Llama2-7b and Mistral-7b) demonstrate reasonable performance on SummEval and TopicalChat, but larger models (GPT3.5) perform much better, and when applying absolute scoring can outperform smaller systems using comparative assessment.
6.2 Attack on Surrogate Model
Section 5.3 details the attack approach to learn the universal attack phrases for the surrogate model. Figure 2 illustrates the impact of the universal adversarial on SummEval and TopicalChat, where FlanT5-xl is used as the surrogate LLM assessment system. For Summeval, the overall score (OVE) and consistency (CON) is attacked while for Topical-Chat the overall score (OVE) and continuity (CNT) is attacked. The attributes CON and CNT were selected due to the similar performance for these attributes in the absolute and comparative settings (seen in Tables 1 and 2).
The success of the adversarial attacks is measured by the average ranks of the text after an attack. Figure 2 demonstrates that both comparative assessment and absolute assessment systems have some vulnerability to adversarial attacks, as the average rank decreases, and continues to decrease as more words are added to the attack phrase. However, absolute scoring systems are significantly more susceptible to universal adversarial attacks, and with just four universal attack words, the absolute scoring system will consistently provide a rank of 1 to nearly all input texts. Table 3 provides the raw scores for comparative and absolute assessment, where we see that for absolute assessment, a universal attack phrase of 4 words will yield assessment scores on average near the maximum score of 5. The specific universal attack phrases learnt for each task are given in Appendix A.
@@ 6.2 Attack on Surrogate Model
Figure 3: Transferability of universal attack phrases from surrogate FlanT5-xl to target models.
The relative robustness of comparative assessment systems over absolute assessment systems can perhaps be explained intuitively. In an absolute assessment setting, an adversary exploits an input space which is not well understood by the model and identifies a region that spuriously encourages the model to predict a high score. However, in comparative assessment, the model is forced to compare the quality of the attacked text to another (unattacked) text, meaning the attack phrase learnt has to be invariant to the text used for comparison. This makes it more challenging to find an effective universal attack phrase. Further explanations for the relative robustness of comparative assessment systems are explored in Appendix B.
6.3 Transferability of the Surrogate Attack
Figure 2 demonstrated that absolute assessment systems are highly vulnerable to a simple universal attack phrase concatenated to an input text. To evaluate the effectiveness of these attack phrases on more powerful target models, we explicitly transfer the attacks learned on the FlanT5-xl surrogate model to other models such as Llama2, Mistral and GPT3.5. We focus on transferring the absolute scoring attacks, as comparative assessments were found to be relatively robust for the surrogate FlanT5-xl model. Figure 3 shows the results of transferring the attack phrases to these models, highlighting several key findings: 1) There can be a high level of attack transferability for absolute scoring. For TopicalChat, the attacks generalize very well to nearly all systems, with all systems being very susceptible to attacks when assessing continuity. 2) When more powerful models assess the overall (OVE) quality, the transferability is less effective, suggesting that assessing more general, abstract qualities can be more robust. Interestingly, powerful large models (GPT3.5) are more susceptible when attacked by shorter phrases, possibly because longer phrases may begin to overfit the properties of the surrogate model. 3) The attack transfers with mixed success for SummEval, which may highlight that the complexity of the dataset can influence attack transferability.
@@ 6.4 Attack Detection
In this section, we perform an initial investigation into possible defences that could be applied to detect if an adversary is exploiting a system. Defences can take two forms: adversarial training (Goodfellow et al., 2015) where the LLM is re-trained with adversarial examples, or adversarial attack detection where a separate module is designed to identify adversarial inputs. Although recent LLM adversarial training approaches have been proposed (Zhou et al., 2024; Zhang et al., 2023b), re-training is computationally expensive and can harm model performance, hence detection is preferred. Recent detection approaches for NLG adversarial attacks tend to focus on attacks that circumvent LLM safety filters, e.g., generating malicious content by jailbreaking (Liu et al., 2023c; Zou et al., 2023; Jin et al., 2024). Robey et al. (2023) propose SmoothLLM, where multiple versions of the perturbed input are passed to an LLM and the outputs aggregated. Such defences are inappropriate for LLM-as-a-judge setups, as though the perturbations are designed to cause no semantic change, they can result in changes in other attributes, such as fluency and style, which will impact the LLM assessment. Similarly, Jain et al. (2023); Kumar et al. (2024) propose defence approaches that involve some form of paraphrasing or filtering of the input sequence, which again interferes with the LLM-as-a-judge scores.
@@
$\texttt{P}=\frac{\texttt{TP}}{\texttt{TP+FP}}\quad\texttt{R}=\frac{\texttt{TP}}{\texttt{TP+FN}}\quad\texttt{F1}=2\cdot\frac{\texttt{P}\cdot\texttt{R}}{\texttt{P}+\texttt{R}},$
@@ 7 Conclusions
This is the first work to examine the adversarial robustness of zero-shot LLM assessment methods against universal adversarial attacks, and reveal significant vulnerabilities in LLM absolute scoring and mild vulnerabilities in LLM comparative assessment. We demonstrate that the same short 4-word universal adversarial can be appended to any input text to deceive LLM assessment system into predicting inflated scores. Notably, LLM-scoring attacks developed with a smaller surrogate LLM-scoring system can be effectively transferred to larger LLMs such as ChatGPT. We also provide an initial investigation into simple detection approaches, and show that perplexity can be a promising tool for identifying adversarially manipulated inputs. Further work can explore adaptive attacks and more sophisticated defence approaches to minimize the risk of misuse. On the whole, this paper raises awareness around the susceptibility of LLM-as-a-judge NLG assessment systems to universal and transferable adversarial attacks.